# Steer-Bench: A Benchmark for Evaluating the Steerability of Large Language Models

### v1

@@ Appendix B Data statistics
1 | ai, features, will, learning, generative, apple, machine, cloud, iphone, it | |
2 | wallet, nfc, payment, tap, digital, contactless, qr, banking, app, debit | |
3 | carplay, tesla, infotainment, automotive, vehicles, android, manufacturers, bmw, toyota, evs | |
5 | airpods, headphones, buds, pair, earbuds, bose, audio, cancellation, sony, beats | |
6 | camera, cameras, processing, quality, photography, sensors, pixel, iphone, telephoto, phones | |
8 | android, iphone, google, phones, apps, os, switch, apple, samsung, back | |
9 | pixel, pixels, samsung, phones, 8a, xl, android, nexus, hardware, camera | |
11 | foldable, folding, foldables, folds, phones, tablet, slab, hinge, unfolded, screens | |
14 | apps, developers, malware, security, android, appstore, alternative, ios, review, apple | |
16 | chrome, browser, safari, firefox, browsers, webkit, mozilla, google, eu, blink | |
17 | imessage, whatsapp, sms, messaging, telegram, messenger, texting, android, texts, iphone | |
18 | comment, opinion, comments, argument, post, downvoted, read, know, thread, troll | |
19 | airtag, network, tags, trackers, pebblebee, bluetooth, google, wallet, locate, ping | |
@@
2 | comment, post, this, people, reddit, what, re, read, know, like | |
7 | windows, win11, w11, win10, upgrade, microsoft, it, w10, ui, eol | |
10 | mac, macos, apple, macbook, linux, macs, os, hardware, laptop, air | |
12 | office, libreoffice, excel, onlyoffice, libre, word, microsoft, openoffice, docs, version | |
13 | laptop, laptops, ram, dell, thinkpad, cpu, thinkpads, intel, hardware, ssd | |
15 | keyboard, mouse, ctrl, shortcuts, gestures, layout, touchpad, shortcut, trackpad, button | |
19 | partition, ssd, drives, partitions, ntfs, hdd, ssds, install, sata, gparted |
@@
1 | insurance, healthcare, medicare, universal, companies, deductible, medicaid, hospitals, premiums, claims | |
2 | hours, overtime, hourly, salaried, salary, shifts, 10, per, schedule, workweek | |
3 | taxes, income, taxed, wealth, taxing, billionaires, government, irs, stocks, wealthy | |
4 | ai, automation, robots, art, jobs, technology, automated, automate, machines, artists | |
5 | mcdonald, burger, wendy, franchise, restaurants, prices, fries, chipotle, burgers, menu | |
6 | interviews, hiring, applications, applicants, resumes, process, interviewing, recruiters, interviewer, tests | |
7 | union, unions, dues, unionize, unionized, benefits, contract, unionizing, company, employees | |
8 | billionaires, wealth, billionaire, billion, wealthy, people, millionaires, world, society, hoarding | |
9 | unemployment, fired, employment, employer, termination, fire, severance, department, claim, attorney | |
10 | ceo, ceos, shareholders, profits, shareholder, companies, business, investors, executives, shares | |
11 | boomers, generation, boomer, millennials, generations, older, millennial, parents, generational, genx | |
12 | she, boss, job, was, manager, tell, herself, it, sounds, has | |
13 | argument, read, wrong, what, comments, conversation, arguing, thread, response, opinion | |
14 | kamala, voters, democrats, bernie, election, dems, democratic, pelosi, dnc, voting | |
15 | pto, unlimited, vacation, hours, accrued, company, policy, week, pay, request | |
16 | office, home, hybrid, work, remotely, from, working, company, productive, commute | |
17 | raise, raises, increase, pay, inflation, job, salary, boss, than, months | |
18 | minimum, wage, 25, federal, wages, inflation, increase, hr, prices, workers | |
19 | break, unpaid, minutes, hours, lunches, clock, shift, min, law, work | |
@@
1 | russia, ukraine, putin, nato, war, ukrainians, nuclear, europe, invaded, zelensky | |
2 | comment, your, argument, facts, point, conversation, wrong, lol, read, evidence | |
3 | phones, data, google, ads, app, 5g, devices, android, privacy, iphone | |
4 | bible, church, christian, christianity, religious, christians, religions, catholic, believe, sin | |
5 | trans, gender, transgender, dysphoria, children, being, their, lgbtq, cis, is | |
6 | conspiracy, conspiracies, sub, theories, theory, theorists, subreddit, theorist, believe, political | |
7 | gun, guns, firearms, amendment, shootings, laws, militia, firearm, regulated, nra | |
8 | police, cops, officer, officers, training, law, protect, shooting, gun, duty | |
9 | assassination, bullet, attempt, shots, head, snipers, shooting, president, blood, the | |
10 | border, immigration, immigrants, migrants, illegals, immigrant, wall, undocumented, republicans, migrant | |
11 | biden, trump, debate, dementia, president, joe, election, democrats, obama, who | |
12 | abortion, abortions, fetus, roe, woman, rights, embryos, states, law, choice | |
13 | insurance, healthcare, companies, ceo, medicare, aca, claims, medicaid, premiums, private | |
14 | drugs, smoking, nicotine, marijuana, tobacco, addiction, meth, vape, cigarettes, opioids | |
15 | war, terrorism, terrorist, civilians, terrorists, ww3, wars, iraq, bombing, world | |
16 | jury, trial, judge, convicted, prosecution, court, defendant, prosecutors, verdict, convict | |
17 | education, schools, teachers, college, degree, colleges, university, universities, district, admissions | |
18 | climate, co2, fossil, emissions, earth, scientists, atmosphere, environmental, are, science | |
19 | parents, children, parent, bullying, bullies, school, bullied, bully, adults, family | |
@@
1 | surgery, vasectomy, procedure, hysterectomy, iud, insurance, sterilization, periods, pill, pregnancy | |
2 | eat, food, eating, foods, meal, meals, eats, fat, cook, protein | |
3 | gap, sibling, siblings, apart, age, each, another, baby, having, sister | |
4 | cat, pet, pets, animal, love, my, have, are, vet, don | |
5 | names, call, nickname, mama, grandma, his, nicknames, change, mr, use | |
6 | party, birthday, parties, cake, birthdays, friends, invites, family, host, attend | |
7 | she, wants, abortion, if, will, it, decision, is, tell, for | |
8 | gifts, gift, christmas, toys, presents, birthday, things, giving, box, items | |
12 | trip, disney, vacation, travel, vacations, beach, fun, hotel, family, park | |
13 | books, read, reading, library, letters, phonics, reader, dyslexia, learning, reads | |
14 | post, op, comments, read, advice, opinion, re, downvoted, troll, thread | |
15 | mil, grandparents, mom, parents, family, relationship, husband, grandparent, grandkids, visit | |
17 | friends, friend, friendship, friendships, new, with, group, kids, hang, she | |
19 | religion, church, religious, christian, catholic, religions, bible, christianity, agnostic, christians |
@@
1 | philosophy, plato, philosophers, read, philosopher, socrates, books, academic, history, science | |
3 | determinism, deterministic, compatibilism, freedom, responsibility, control, libertarian, choose, universe, causal | |
4 | conservatives, election, democrats, liberal, republican, politics, democratic, conservatism, government, sides | |
5 | racism, race, crime, racial, poverty, whites, cops, races, gangs, immigrants | |
9 | consciousness, brain, physicalism, qualia, idealism, identity, dualism, self, states, body | |
10 | comment, response, point, didn, understand, sorry, appreciate, op, post, this | |
11 | abortion, abortions, fetus, rights, roe, argument, babies, antinatalism, autonomy, procreation | |
@@
11 | trans, gender, transgender, dysphoria, identity, cis, puberty, intersex, surgery, transitioning | |
12 | your, comment, re, argument, point, discussion, responding, read, understand, sorry | |
13 | meat, vegan, vegetarian, vegans, veganism, diets, animals, vegetarians, foods, chicken | |
15 | ai, human, intelligence, data, it, machine, robots, robot, agi, learning | |
@@
1 | climate, change, science, is, scientists, about, people, what, warming, believe | |
2 | meat, cows, eat, vegan, animal, beef, methane, agriculture, farmers, cattle | |
3 | solar, wind, panels, renewables, electricity, coal, are, turbines, renewable, but | |
4 | florida, rise, insurance, beach, tide, will, change, climate, property, flood | |
5 | she, to, kamala, trump, is, taylor, fracking, vote, if, climate | |
8 | heat, summer, temperatures, degrees, weather, winter, year, days, record, deaths | |
10 | china, emissions, us, capita, world, coal, countries, are, india, ccp | |
11 | oil, fossil, fuels, companies, industry, big, are, gas, exxon, money | |
12 | your, what, comment, troll, don, are, question, have, read, know | |
13 | science, scientific, scientists, the, study, consensus, peer, what, studies, research | |
16 | nuclear, reactors, waste, energy, solar, plants, build, wind, are, more | |
@@
3 | calories, macros, app, tracking, cronometer, carbs, calculator, per, weigh, body | |
4 | b12, supplements, vitamins, multivitamin, supplementation, yeast, deficiencies, supplementing | |
5 | bring, birthday, family, party, restaurant, dinner, vegan, make, invited, guests | |
7 | ibs, constipation, psyllium, bloating, digestive, stool, probiotics, microbiome, crohn, bacteria | |
12 | cholesterol, ldl, statins, triglycerides, lipid, keto, arteries, elevated, fats, dr | |
15 | thank, appreciate, journey, yourself, sharing, advice, proud, happy, congratulations, helpful | |
@@
2 | air, pan, cook, fryer, steak, sear, oven, sous, iron, cooking | |
4 | milk, dairy, cream, cheese, yogurt, oat, butter, kefir, coffee, raw | |
@@
5 | math, calculator, memorization, fractions, division, teach, basic, memorize, calculators, tables | |
7 | religion, bible, religious, church, commandments, beliefs, religions, christianity, muslim, atheist | |
8 | reading, read, phonics, dyslexia, letters, spelling, dyslexic, word, program, blending | |
9 | post, comment, comments, op, what, sub, response, opinion, point, read | |
14 | ai, using, writing, write, generated, essay, students, can, prompt, plagiarism | |
16 | parents, children, parenting, parent, people, trauma, adults, think, be, life | |
18 | summer, week, year, august, june, thanksgiving, school, february, month, during |
@@
1 | inflation, cut, recession, economy, cpi, data, gdp, jobs, banks, government | |
2 | tesla, elon, musk, ev, vehicles, robotaxi, teslas, waymo, uber, charging | |
3 | cars, vehicle, maintenance, toyota, financing, buy, honda, dealership, cash, engine | |
4 | comment, post, this, sub, regarded, advice, appreciate, do, like, sorry | |
7 | job, degree, jobs, college, career, hours, salary, field, pay, skills | |
10 | taxes, irs, withholding, refund, 1099, filing, w2, deductions, cpa, married | |
@@
5 | winter, heating, heater, cabin, battery, ac, weather, temperatures, preconditioning, climate | |
8 | pedal, regen, braking, brakes, accelerator, friction, speed, control, pads, manual | |
10 | trump, biden, democrats, party, election, speech, gop, voters, politics, media | |
16 | truck, trucks, towing, pickups, beds, hauling, cabs, hitch, payload, big | |
19 | cybertruck, ct, truck, cybertrucks, trucks, tesla, r1t, towing, silverado, vehicle | |
@@
5 | comment, post, joke, read, downvoted, re, word, sorry, grammar, argument | |
7 | cult, mormon, religion, scientology, churches, mormons, missionaries, jehovah, cults, satanic | |
11 | thank, sharing, appreciate, wish, love, post, beautiful, words, journey, happy | |
15 | clothes, wardrobe, wool, clothing, thrift, brands, outfits, stores, thrifting, items | |
17 | ads, spotify, plex, vpn, hulu, android, subscriptions, torrent, services, piracy | |
@@
1 | electronic, album, house, techno, and, stuff, from, albums, music, check | |
3 | spotify, cds, streaming, classical, music, artists, it, vinyl, app, radio | |
@@
17 | her, to, it, with, but, can, will, know, support, baby |

### v2

@@ Appendix B Data statistics
| 1 | ai, features, will, learning, generative, apple, machine, cloud, iphone, it | |
| 2 | wallet, nfc, payment, tap, digital, contactless, qr, banking, app, debit | |
| 3 | carplay, tesla, infotainment, automotive, vehicles, android, manufacturers, bmw, toyota, evs | |
| 5 | airpods, headphones, buds, pair, earbuds, bose, audio, cancellation, sony, beats | |
| 6 | camera, cameras, processing, quality, photography, sensors, pixel, iphone, telephoto, phones | |
| 8 | android, iphone, google, phones, apps, os, switch, apple, samsung, back | |
| 9 | pixel, pixels, samsung, phones, 8a, xl, android, nexus, hardware, camera | |
| 11 | foldable, folding, foldables, folds, phones, tablet, slab, hinge, unfolded, screens | |
| 14 | apps, developers, malware, security, android, appstore, alternative, ios, review, apple | |
| 16 | chrome, browser, safari, firefox, browsers, webkit, mozilla, google, eu, blink | |
| 17 | imessage, whatsapp, sms, messaging, telegram, messenger, texting, android, texts, iphone | |
| 18 | comment, opinion, comments, argument, post, downvoted, read, know, thread, troll | |
| 19 | airtag, network, tags, trackers, pebblebee, bluetooth, google, wallet, locate, ping | |
@@
| 2 | comment, post, this, people, reddit, what, re, read, know, like | |
| 7 | windows, win11, w11, win10, upgrade, microsoft, it, w10, ui, eol | |
| 10 | mac, macos, apple, macbook, linux, macs, os, hardware, laptop, air | |
| 12 | office, libreoffice, excel, onlyoffice, libre, word, microsoft, openoffice, docs, version | |
| 13 | laptop, laptops, ram, dell, thinkpad, cpu, thinkpads, intel, hardware, ssd | |
| 15 | keyboard, mouse, ctrl, shortcuts, gestures, layout, touchpad, shortcut, trackpad, button | |
| 19 | partition, ssd, drives, partitions, ntfs, hdd, ssds, install, sata, gparted |
@@
| 1 | insurance, healthcare, medicare, universal, companies, deductible, medicaid, hospitals, premiums, claims | |
| 2 | hours, overtime, hourly, salaried, salary, shifts, 10, per, schedule, workweek | |
| 3 | taxes, income, taxed, wealth, taxing, billionaires, government, irs, stocks, wealthy | |
| 4 | ai, automation, robots, art, jobs, technology, automated, automate, machines, artists | |
| 5 | mcdonald, burger, wendy, franchise, restaurants, prices, fries, chipotle, burgers, menu | |
| 6 | interviews, hiring, applications, applicants, resumes, process, interviewing, recruiters, interviewer, tests | |
| 7 | union, unions, dues, unionize, unionized, benefits, contract, unionizing, company, employees | |
| 8 | billionaires, wealth, billionaire, billion, wealthy, people, millionaires, world, society, hoarding | |
| 9 | unemployment, fired, employment, employer, termination, fire, severance, department, claim, attorney | |
| 10 | ceo, ceos, shareholders, profits, shareholder, companies, business, investors, executives, shares | |
| 11 | boomers, generation, boomer, millennials, generations, older, millennial, parents, generational, genx | |
| 12 | she, boss, job, was, manager, tell, herself, it, sounds, has | |
| 13 | argument, read, wrong, what, comments, conversation, arguing, thread, response, opinion | |
| 14 | kamala, voters, democrats, bernie, election, dems, democratic, pelosi, dnc, voting | |
| 15 | pto, unlimited, vacation, hours, accrued, company, policy, week, pay, request | |
| 16 | office, home, hybrid, work, remotely, from, working, company, productive, commute | |
| 17 | raise, raises, increase, pay, inflation, job, salary, boss, than, months | |
| 18 | minimum, wage, 25, federal, wages, inflation, increase, hr, prices, workers | |
| 19 | break, unpaid, minutes, hours, lunches, clock, shift, min, law, work | |
@@
| 1 | russia, ukraine, putin, nato, war, ukrainians, nuclear, europe, invaded, zelensky | |
| 2 | comment, your, argument, facts, point, conversation, wrong, lol, read, evidence | |
| 3 | phones, data, google, ads, app, 5g, devices, android, privacy, iphone | |
| 4 | bible, church, christian, christianity, religious, christians, religions, catholic, believe, sin | |
| 5 | trans, gender, transgender, dysphoria, children, being, their, lgbtq, cis, is | |
| 6 | conspiracy, conspiracies, sub, theories, theory, theorists, subreddit, theorist, believe, political | |
| 7 | gun, guns, firearms, amendment, shootings, laws, militia, firearm, regulated, nra | |
| 8 | police, cops, officer, officers, training, law, protect, shooting, gun, duty | |
| 9 | assassination, bullet, attempt, shots, head, snipers, shooting, president, blood, the | |
| 10 | border, immigration, immigrants, migrants, illegals, immigrant, wall, undocumented, republicans, migrant | |
| 11 | biden, trump, debate, dementia, president, joe, election, democrats, obama, who | |
| 12 | abortion, abortions, fetus, roe, woman, rights, embryos, states, law, choice | |
| 13 | insurance, healthcare, companies, ceo, medicare, aca, claims, medicaid, premiums, private | |
| 14 | drugs, smoking, nicotine, marijuana, tobacco, addiction, meth, vape, cigarettes, opioids | |
| 15 | war, terrorism, terrorist, civilians, terrorists, ww3, wars, iraq, bombing, world | |
| 16 | jury, trial, judge, convicted, prosecution, court, defendant, prosecutors, verdict, convict | |
| 17 | education, schools, teachers, college, degree, colleges, university, universities, district, admissions | |
| 18 | climate, co2, fossil, emissions, earth, scientists, atmosphere, environmental, are, science | |
| 19 | parents, children, parent, bullying, bullies, school, bullied, bully, adults, family | |
@@
| 1 | surgery, vasectomy, procedure, hysterectomy, iud, insurance, sterilization, periods, pill, pregnancy | |
| 2 | eat, food, eating, foods, meal, meals, eats, fat, cook, protein | |
| 3 | gap, sibling, siblings, apart, age, each, another, baby, having, sister | |
| 4 | cat, pet, pets, animal, love, my, have, are, vet, don | |
| 5 | names, call, nickname, mama, grandma, his, nicknames, change, mr, use | |
| 6 | party, birthday, parties, cake, birthdays, friends, invites, family, host, attend | |
| 7 | she, wants, abortion, if, will, it, decision, is, tell, for | |
| 8 | gifts, gift, christmas, toys, presents, birthday, things, giving, box, items | |
| 12 | trip, disney, vacation, travel, vacations, beach, fun, hotel, family, park | |
| 13 | books, read, reading, library, letters, phonics, reader, dyslexia, learning, reads | |
| 14 | post, op, comments, read, advice, opinion, re, downvoted, troll, thread | |
| 15 | mil, grandparents, mom, parents, family, relationship, husband, grandparent, grandkids, visit | |
| 17 | friends, friend, friendship, friendships, new, with, group, kids, hang, she | |
| 19 | religion, church, religious, christian, catholic, religions, bible, christianity, agnostic, christians |
@@
| 1 | philosophy, plato, philosophers, read, philosopher, socrates, books, academic, history, science | |
| 3 | determinism, deterministic, compatibilism, freedom, responsibility, control, libertarian, choose, universe, causal | |
| 4 | conservatives, election, democrats, liberal, republican, politics, democratic, conservatism, government, sides | |
| 5 | racism, race, crime, racial, poverty, whites, cops, races, gangs, immigrants | |
| 9 | consciousness, brain, physicalism, qualia, idealism, identity, dualism, self, states, body | |
| 10 | comment, response, point, didn, understand, sorry, appreciate, op, post, this | |
| 11 | abortion, abortions, fetus, rights, roe, argument, babies, antinatalism, autonomy, procreation | |
@@
| 11 | trans, gender, transgender, dysphoria, identity, cis, puberty, intersex, surgery, transitioning | |
| 12 | your, comment, re, argument, point, discussion, responding, read, understand, sorry | |
| 13 | meat, vegan, vegetarian, vegans, veganism, diets, animals, vegetarians, foods, chicken | |
| 15 | ai, human, intelligence, data, it, machine, robots, robot, agi, learning | |
@@
| 1 | climate, change, science, is, scientists, about, people, what, warming, believe | |
| 2 | meat, cows, eat, vegan, animal, beef, methane, agriculture, farmers, cattle | |
| 3 | solar, wind, panels, renewables, electricity, coal, are, turbines, renewable, but | |
| 4 | florida, rise, insurance, beach, tide, will, change, climate, property, flood | |
| 5 | she, to, kamala, trump, is, taylor, fracking, vote, if, climate | |
| 8 | heat, summer, temperatures, degrees, weather, winter, year, days, record, deaths | |
| 10 | china, emissions, us, capita, world, coal, countries, are, india, ccp | |
| 11 | oil, fossil, fuels, companies, industry, big, are, gas, exxon, money | |
| 12 | your, what, comment, troll, don, are, question, have, read, know | |
| 13 | science, scientific, scientists, the, study, consensus, peer, what, studies, research | |
| 16 | nuclear, reactors, waste, energy, solar, plants, build, wind, are, more | |
@@
| 3 | calories, macros, app, tracking, cronometer, carbs, calculator, per, weigh, body | |
| 4 | b12, supplements, vitamins, multivitamin, supplementation, yeast, deficiencies, supplementing | |
| 5 | bring, birthday, family, party, restaurant, dinner, vegan, make, invited, guests | |
| 7 | ibs, constipation, psyllium, bloating, digestive, stool, probiotics, microbiome, crohn, bacteria | |
| 12 | cholesterol, ldl, statins, triglycerides, lipid, keto, arteries, elevated, fats, dr | |
| 15 | thank, appreciate, journey, yourself, sharing, advice, proud, happy, congratulations, helpful | |
@@
| 2 | air, pan, cook, fryer, steak, sear, oven, sous, iron, cooking | |
| 4 | milk, dairy, cream, cheese, yogurt, oat, butter, kefir, coffee, raw | |
@@
| 5 | math, calculator, memorization, fractions, division, teach, basic, memorize, calculators, tables | |
| 7 | religion, bible, religious, church, commandments, beliefs, religions, christianity, muslim, atheist | |
| 8 | reading, read, phonics, dyslexia, letters, spelling, dyslexic, word, program, blending | |
| 9 | post, comment, comments, op, what, sub, response, opinion, point, read | |
| 14 | ai, using, writing, write, generated, essay, students, can, prompt, plagiarism | |
| 16 | parents, children, parenting, parent, people, trauma, adults, think, be, life | |
| 18 | summer, week, year, august, june, thanksgiving, school, february, month, during |
@@
| 1 | inflation, cut, recession, economy, cpi, data, gdp, jobs, banks, government | |
| 2 | tesla, elon, musk, ev, vehicles, robotaxi, teslas, waymo, uber, charging | |
| 3 | cars, vehicle, maintenance, toyota, financing, buy, honda, dealership, cash, engine | |
| 4 | comment, post, this, sub, regarded, advice, appreciate, do, like, sorry | |
| 7 | job, degree, jobs, college, career, hours, salary, field, pay, skills | |
| 10 | taxes, irs, withholding, refund, 1099, filing, w2, deductions, cpa, married | |
@@
| 5 | winter, heating, heater, cabin, battery, ac, weather, temperatures, preconditioning, climate | |
| 8 | pedal, regen, braking, brakes, accelerator, friction, speed, control, pads, manual | |
| 10 | trump, biden, democrats, party, election, speech, gop, voters, politics, media | |
| 16 | truck, trucks, towing, pickups, beds, hauling, cabs, hitch, payload, big | |
| 19 | cybertruck, ct, truck, cybertrucks, trucks, tesla, r1t, towing, silverado, vehicle | |
@@
| 5 | comment, post, joke, read, downvoted, re, word, sorry, grammar, argument | |
| 7 | cult, mormon, religion, scientology, churches, mormons, missionaries, jehovah, cults, satanic | |
| 11 | thank, sharing, appreciate, wish, love, post, beautiful, words, journey, happy | |
| 15 | clothes, wardrobe, wool, clothing, thrift, brands, outfits, stores, thrifting, items | |
| 17 | ads, spotify, plex, vpn, hulu, android, subscriptions, torrent, services, piracy | |
@@
| 1 | electronic, album, house, techno, and, stuff, from, albums, music, check | |
| 3 | spotify, cds, streaming, classical, music, artists, it, vinyl, app, radio | |
@@
| 17 | her, to, it, with, but, can, will, know, support, baby |